# A quantitative analysis of semantic information in deep representations of text and images

### v3

@@ Appendix I Binary Intrinsic Dimension of DeepSeek-V3 representations
Results are in Fig. 13. In all cases we compute the information imbalance (II) in both directions (from the source model to the target and vice-versa), always indicating the direction in the figure legends. The DeepSeek-V3 $\to$ CLIP-Image comparison directly tests whether a visual encoder trained to align with text predicts textual representations better than a unimodal vision model. Its minimum II is approximately 0.3 ($\approx 15\%$ shared features), which is lower than the $\approx 0.40$ minimum for DeepSeek-V3 $\to$ DinoV2 (main text Fig. 4). The DinoV2 $\to$ CLIP-Image II remains extremely low ($\approx 0.01$) past a relative depth of $\approx 0.8$, consistent with a near-identical visual representation. Conversely, the DeepSeek-V3 $\to$ CLIP-Text II decreases more gradually and stabilizes around 0.4, indicating weaker alignment of CLIP’s text branch to purely textual models. Shaded regions in Fig. 13 show 95% bootstrap confidence intervals.
@@
Figure 13: Information imbalance (II) on flickr30k involving CLIP representations. Top: DinoV2 image vs. CLIP image. Middle: DeepSeek-V3 text (layer 52) vs. CLIP text. Bottom: DeepSeek-V3 text vs. CLIP image (new cross-modal analysis). Lower II indicates stronger semantic predictivity. Shaded areas show 95% bootstrap confidence intervals.

### v4

@@ Appendix I Token-Level Information Imbalance and Causal Asymmetry Across Languages
Results are in Fig. 14. In all cases we compute the Information Imbalance in both directions (from the source model to the target and vice-versa), always indicating the direction in the figure legends. The DeepSeek-V3 $\to$ CLIP-Image comparison directly tests whether a visual encoder trained to align with text predicts textual representations better than a unimodal vision model. Its minimum Information Imbalance is approximately 0.3 ($\approx 15\%$ shared features), which is lower than the $\approx 0.40$ minimum for DeepSeek-V3 $\to$ DinoV2 (main text Fig. 4). The DinoV2 $\to$ CLIP-Image Information Imbalance remains extremely low ($\approx 0.01$) past a relative depth of $\approx 0.8$, consistent with a near-identical visual representation. Conversely, the DeepSeek-V3 $\to$ CLIP-Text Information Imbalance decreases more gradually and stabilizes around 0.4, indicating weaker alignment of CLIP’s text branch to purely textual models. Shaded regions in Fig. 14 show 95% bootstrap confidence intervals.
@@
Figure 14: Information Imbalance on flickr30k involving CLIP representations. Top: DinoV2 image vs. CLIP image. Middle: DeepSeek-V3 text (layer 52) vs. CLIP text. Bottom: DeepSeek-V3 text vs. CLIP image (new cross-modal analysis). Lower Information Imbalance indicates stronger semantic predictivity. Shaded areas show 95% bootstrap confidence intervals.